Jointsoft is a great over-the-counter arthritis medication, but who will ever know about it? Unfortunately, many people with arthritis tend to be elderly and rather immobile, so advertisers of arthritis medications face limitations in ways to get their messages across. Currently, their best modes of advertisement are commercials on daytime TV, advertisements in select magazines, fliers in convalescent homes, and (believe it or not) advertisements on certain Web pages Marketing managers for Jointsoft are investigating these four modes of advertisement in four small communities (with a different mode of advertisement in each community). The marketing managers have selected 39 days at random and are looking at the daily sales (in dollars) in each of the communities on each of these days. Here is what they have to work with: Sample Sample Sample size Groups variance 2224.6 1743.9 1991.6 3334.2 mean TV Magazines Fliers Web pages 39 571.2 590.3 559.2 580.6 39 39 Epica Send data to Excel Suppose that the marketing managers perform a one-way, independent-samples ANOVA test to decide if there are differences in the mean daily sales arising from the four modes of advertisement. (So, theyre assuming that the only difference among the four communities is the mode of advertisement used in it.) Such a test uses the statistic Variation between the samples Variation within the samples For the information in the chart above, F 2.96
Answer #1

Ans:

Given that

F=2.96

number of subjects=4*39=156

nuber of groups=4

df(numerator)=4-1=3

df(denominator)=156-4=152

p-value=FDIST(2.96,3,152)=0.0342

As,p-value>0.01,we fail to reject null hypothesis

Marketing managers can not conclude that mean sales arising from atleast one of the modes of advertisement differs from others.